According to the Bohrs model of hydrogen atom te electric current generated due to the motion of the electron in nth orbit is given by?

Grade:

1 Answers

vikas askiitian expert
13 years ago

radius (rn) = n2ao/Z  .........1            (n=orbit , Z = atomic no)

  ao = 52.9pm

velocity(Vn) = VoZ/n       ..........2                                  (Vo = 2.18*106m/s)

time period of revolution = 2pir/V

                              Tn   =  2pin3ao/Z2Vo

now , in hydrogen atom ,electron moves in circular orbit , electron is charge particle & due to motion of any charge particle electric current is gererated so

  I = dq/dt

  I = e/Tn = Z2Voe/2piaon3                         (e is electronic charge)

this is the required expression  .......

approve my ans if u like it